An alien from the planet OMO Centauri writes the first ten prime numbers in arbitrary order as U,W, XW, ZZ, V, Y, ZV, ZW, ZY, and X. Each letter represents a nonzero digit. Each letter represents the same digit everywhere it appears, and different letters represent different digits. Also, the alien is using a base other than base ten. The alien writes another number as UZWX. Compute this number (expressed in base ten, with the usual, human digits).
1
Expert's answer
2020-12-15T01:08:58-0500

First 5 prime numbers is 2,3,5,7,11

sixth prime number is 13

We have 5 numbers which have one letter

So Base could be only 12 - between 11 and 13


Next, we have 4 prime numbers with first letter Z

ZZ - should be 1*12 + 1 = 13


ZV, ZW, ZY is the next 3 prime numbers 17,19,23

17 = Z + 5

19 = Z + 7

23 = Z + 11

Hence V,W,Y - represents 5,7,11 in any order, the rest letter U is 3


Next, the last number XW - is 29

X*12+W = 2*24 + W = 29

Hence X is 2 and W is 5

So Z = 1, X = 2,U = 3 and W = 5


Answer: UZWX = 3152